# Lookahead: An Inference Acceleration Framework for Large Language Model with Lossless Generation Accuracy

### v2

@@ 4.2. Multi-Branch Draft
Fortunately, through careful observation of the context, we have noticed that certain branches in the drafts have common prefix tokens. For example, in Figure 2, the branches [on, my, knee] and [on, a, table] both share the prefix token [on]. By leveraging this observation, we can compress and organize the multi-branch draft using a hierarchical structure, allowing for the inclusion of additional branches while maintaining the same token capacity. To achieve this, we recursively merge the prefix token(s) shared by multiple branches, and then append the remaining tokens of each branch. We refer to these compressed and organized drafts as hierarchical multi-branch draft. On the other hand, drafts organized in a parallel manner without using the hierarchical structure are referred to as parallel multi-branch draft.
As illustrated in Figure 2 describing the process of an inference step, through the single-branch strategy, the single draft [on, my, knee] is appended to the next token [sits] directly for the single forward process in VA, which in turn gives the output [on, a]. Being benefited from the multi-branch strategy with the parallel multi-branch draft, two branches [on, my, knee] and [on, a, table] are organized in parallel and appended to the next token [sits] for verifying, which outputs [on, a, chair]. By utilizing the hierarchical multi-branch draft, the shared prefix token [on] are merged, after which the token list [my, knee, a, table] are appended. By doing so, given the same decoding length, we are able to save an additional space to append one more token [chair] from another branch [on, a, chair]. Combined with the next token, we collect the token list [sits, on, my, knee, a, table, chair] for decoding through the single forward process and output [on, a, chair, and]. Therefore, compared to the single-branch strategy, our multi-branch strategy, particularly when employing the hierarchical multi-branch draft, offers the advantage of retrieving multiple drafts, resulting in improved EDL from a statistical perspective and thus significantly enhancing LLMs’ inference speed.
The position IDs and causal masks in a transformer-based model [31] are also merged to align with the merged token list. By doing so, we are able to accommodate more branches compared to the utilization of parallel multi-branch draft, whose details are presented in Figure 3.
To facilitate the hierarchical structure of branches, we utilize a trie tree, a data structure commonly used for efficient retrieval and storage purposes to record more candidate branches, whose details will be introduced in the next section.
@@
4.3. Trie Tree
The trie tree is a data structure that efficiently handles prefix matching by organizing nodes as individual characters or words. In our work, each node of the trie tree represents a token ID, and a path from the root to a leaf node represents a branch.
@@ 4.3.1. Trie Tree Updating
We apply the branch inserting, eliminating and node pruning to update the trie tree, in order to maintain the trie tree’s size and keep the most frequency tokens.
Branch Inserting. Considering that in several typical scenarios (e.g., RAG, Summary), the output are often derived from the prompt, we thus manipulate the prompt into the branches and insert them into the trie tree. Moreover, towards various prompts, there may be similar outputs, e.g., ”keep exercising” in answering ”how to keep healthy” and ”how to be stronger”. Therefore, we also insert the outputs into the trie tree. Given the prompt or the output, each sub-sequence of the successive tokens with the specified length can be regarded as a branch draft and this specified length is the branch length. Each branch is then automatically inserted into the trie tree, whose construction process has been widely applied in document matching [32, 33].
@@
Node Pruning. To maintain the trie tree within a moderate size, in case that the oversized trie tree results in high memory consumption and slow retrieval performance, we develop a pruning strategy. To achieve this, we remove the least frequent nodes dynamically, when the trie tree exceeds a predetermined capacity. By doing so, we are able to optimize memory consumption and improve retrieval performance in our framework.
@@ 4.3.2. Trie Tree Retrieving
We can extract multiple branches from a constructed trie tree by providing a prefix, which is essentially a list of tokens. The length of the prefix influences the number and relevance of the retrieved branches. Short prefixes yield a greater number of branches, while longer prefixes result in more related branches. To strike a balance between the count and correlation of the branches, we adopt a multi-stage retrieval strategy inspired by [16]. Specifically, we begin by attempting to match a longer prefix. If the number of tokens associated with the matched branches is significantly smaller than the CDL, we reduce the length of the prefix and retry the matching process until we obtain a substantial number of tokens linked to matched branches.
If the count of the matched branches falls below a predefined threshold, we utilize all of them for the VA process. However, if the count exceeds a given size, we choose the tokens with the highest frequency. To prioritize the importance of branches derived from the input prompt, we amplify their frequency by a significant factor.
4.4. Work Flow
@@
5. Experiments
5.1. Experimental Setup
We conduct a solid experiment to evaluate the efficacy, robustness and practicality of our Lookahead framework in accelerating LLMs’ inference. Our inference environment includes a server with 32-core CPU, 64 GiB host memory, and a A100-SXM(80G) GPU. Considering the actual industry scenarios that Lookahead is applied, AntRAG is chosen as the evaluation dataset. AntRAG dataset is an internal benchmark dataset that is meticulously collected from a real-life product system in Alipay. The dataset collection process involves submitting user queries to a search engine, which then retrieves the most relevant document alongside the original query. These query-document pairs are combined to form prompts. To ensure the dataset’s quality and coherence, we employ a rigorous process of removing any duplicate or unrelated prompts, resulting in a refined and comprehensive dataset. Furthermore, to validate the robustness and practicality of Lookahead in open-domain scenarios, we additionally conduct experiments utilizing the Dolly dataset  [^3] Dolly dataset: https://huggingface.co/datasets/databricks/databricks-dolly-15k. For the evaluation purposes, considering that all the candidate methods promise the accuracy lossless generation, therefore the inference speed with the unit of output tokens per second is regarded as the unique evaluation metric.
In evaluating the applicability of base LLMs across different datasets, it is essential to consider their adaptability. For our study, we select Antglm-10B as the base model for the AntRAG dataset, while Llama2-13B serves as the base model for the Dolly dataset. Detailed information regarding the datasets and base models utilized can be found in Table 2 and Table 3, respectively. It should be noted that for each dataset, we use the test set containing 1000 samples for the inference speed evaluation, while applying the dev set for warm-up to simulate the real-life scenarios.
@@ 5.1. Experimental Setup
Dataset Split #samples #tokens of prompt #tokens of output
@@
Model Antglm-10B Llama2-13B
Params 10.14B 12.71B
Vocab size 115328 32000
Layer 48 40
Hidden size 4096 5120
Attention head 32 40
MLP intermediate size 16384 13824
5.2. Results
5.2.1. Lookahead’s superior performance
Table 4 exhibits the inference speeds achieved by different acceleration methods. To minimize the influence of random factors, we conduct the evaluation process five times, ensuring the accuracy and reliability. The mean value of the inference speed is then selected as the performance indicator for each acceleration method.
Table 4. Inference Speed of different methods and datasets. The results are calculated by conducting five repeated runs.
Methods AntRAG Dolly
token/s speedup token/s speedup
Baseline 52.4±0.2 1.00 34.0±0.3 1.00
LLMA 165.4±3.3 3.16 50.8±0.2 1.49
Lookahead (Parallel) 263.4±3.3 5.03 68.9±0.4 2.03
Lookahead (Hierarchical) 280.9±4.4 5.36 71.7±0.9 2.11
As depicted in Table 4, the average inference speed of AntGLM-10B towards AntRAG is recorded at 52.4 tokens/s. However, with the incorporation of LLMA, this speed is elevated to 165.4 tokens/s, resulting in a notable 3.16 times speed improvement. Implementing Lookahead with parallel multi-branch draft further enhances AntGLM-10B’s average inference speed to 263.4 tokens/s, achieving a remarkable 5.03 times speed-up, representing a significant acceleration. Additionally, the adoption of hierarchical multi-branch draft via trie tree propels the average inference speed even further, culminating in a 5.36 times speed-up, surpassing LLMA by 70% in terms of acceleration.
In the utilization of Llama2-13B as the base model for inferring the Dolly dataset, the observed speed enhancements resulting from different accelerations may not exhibit significant disparities. It should be noted that the Dolly is an open-domain dataset. In this dataset, the output tokens may not occur in the input prompts, which differs from the real-world scenarios where the output tokens often share a common topic and can be predicted more easily. Nevertheless, Lookahead consistently demonstrates its remarkable superiority and applicability across diverse datasets, further emphasizing its practicality in real-world scenarios.
5.2.2. Hyper-parameters in multi-branch draft
5.2.3. Hyper-parameters in trie tree updating
Table 5. The inference speed of different updating procedures was evaluated, considering various conditions. In the evaluation, ”W/o prompt” refers to the absence of branching from the prompt to the trie tree, while ”W/o output” indicates the absence of branching from the output to the trie tree, ”W/o Pruning” signifies the exclusion of node count reduction in an oversized trie tree through pruning, ”W/o eliminating” denotes the retention of branches derived from the prompt before processing the subsequent query.
Condition W/o prompt W/o output W/o pruning W/o elimin Lookahead
Token/s 234.5 202.0 221.4 234.2 280.9
The capacity of the trie tree has a significant impact on the acceleration performance. To determine the optimal capacity, we conduct empirical experiments, as outlined in Table 6. Instead of using a fixed value, we normalize the node capacity in relation to the decoding length, which allows us to identify a suitable hyper-parameter that performs well across different decoding lengths. Based on our findings, we conclude that node capacities of 16 and 32 times the decoding length yield the best results. Consequently, we have set the default capacity in our framework to be 16 times the decoding length.
Table 6. Inference Speed of different capacity of trie tree. n*DL denotes the capicity is n times the decoding length.
Capacity 1*DL 2*DL 4*DL 8*DL 16*DL 32*DL 64*DL 128*DL
Token/s 254.5 258.7 268.8 273.3 280.7 280.8 279.5 278.3
5.2.4. Lookahead’s efficient memory usage
Figure 6 shows the peak GPU memory consumption for various decoding lengths (1, 2, 4, 8, 16, 32, 64, 128). The decoding length of 1 represents the experience where no Lookahead is used, resulting in a GPU memory consumption of 20.248 GiB. Surprisingly, even with a decoding length of 128, the use of Lookahead only leads to a negligible increase in GPU memory consumption of 0.6%. This minimal increase in memory usage is practically negligible and has no significant impact on real-world applications. Furthermore, we have also evaluated the CPU memory consumption of the trie tree. With the AntRAG dataset, the trie tree only utilizes a mere 260 MiB of memory. This value is negligible when compared to the total memory capacity of a mainstream server.
Figure 6. GPU Memory (i.e., device memory) of different decoding length. The GPU memory denotes the peak memory consumption in a inference process.
5.3. Online Deployment
Our framework has been widely used in real-world scenarios of Alipay since April 2023, due to its accessibility, accuracy-lossless and remarkable acceleration. To be concise, we only report 5 typical scenarios in Table 7. CHART2JSON is a scenario that uses a multi-modal model to convert a chart image to structured content with JSON format, it can achieve extraordinary acceleration due to plenty of template tokens in generated JSONs. Citizen biz agent, enterprise info QA and health suggestion are RAG scenarios, they are used for answering user questions with reference documents. Medical report summary is a scenario to summarize the content (texts after OCR) of a medical report image. With the assistance of our lookahead framework, the latency of all scenarios is decreased significantly while the generation results are the same as the original step-by-step strategy.
Scenario Baseline Lookahead Speedup
CHART2JSON 12.97 2.07 6.26x
Citizen Biz Agent 1.67 0.32 5.21x
Enterprise Info QA 14.01 2.74 5.11x
Health Suggestion 12.41 2.66 4.66x
Medical Report Summary 3.33 1.25 2.66x
6. Conclusion
Appendix A Batch Inference
It is important to note that the aforementioned experiments were conducted with a batch size of 1. However, as illustrated in Table 8, Lookahead also demonstrates its effectiveness with non-singular batch sizes, achieving significantly higher inference speeds compared to step-by-step batch inference and LLMA-based batch inference strategies. Although the speedup is not as remarkable as with singular batch sizes, there are two key reasons for this. Firstly, as the batch size increases, the need for parallel inference begins to strain the available redundancy of the GPU. This can limit the potential speedup that can be achieved. Secondly, different samples within a batch are at different stages of progress, meaning that some may have generated fewer tokens compared to others. This variability in progress leads to additional time consumption in attention calculation, further impacting the overall inference speed. However, it is important to highlight that Lookahead is primarily focused on time-sensitive scenarios rather than throughput-sensitive scenarios. In the future, we plan to optimize the throughput by implementing continuous batching and paged attention mechanisms [34].
Method Batch size=2 Batch size=4
token/s speedup token/s speedup
Baseline 68.0 1.00x 88.2 1.00x
LLMA 185.2 2.72x 214.0 2.43x
Lookahead(Parallel) 274.4 4.04x 291.2 3.30x
Lookahead(Hierarchical) 285.5 4.20x 299.5 3.40x
Appendix B Prompt templates with the Dolly dataset and Llama-13B model.
@@ Appendix A Batch Inference
Similar to [35], we construct a prompt with the below template.

### v3

@@ 4.2. Multi-Branch Draft
4.2.1. Parallel Multi-Branch Draft
The drafts organized in a simple and straight forward parallel manner are referred to as parallel multi-branch draft, whose details can be found in Figure 2 and 3.
4.2.2. Hierarchical Multi-Branch Draft
Fortunately, through careful observation of the context, we have noticed that certain branches in the drafts have common prefix tokens. For example, in Figure 2, the branches [on, my, knee] and [on, a, table] both share the prefix token [on]. By leveraging this observation, we can compress and organize the multi-branch draft using a hierarchical structure, allowing for the inclusion of additional branches while maintaining the same token capacity. To achieve this, we recursively merge the prefix token(s) shared by multiple branches, and then append the remaining tokens of each branch. We refer to these compressed and organized drafts as hierarchical multi-branch draft.
As illustrated in Figure 2 describing the process of an inference step, through the single-branch strategy, the single draft [on, my, knee] is appended to the next token [sits] directly for the single forward process in VA, which in turn gives the output [on, a]. Being benefited from the parallel multi-branch draft, two branches [on, my, knee] and [on, a, table] are organized in parallel and appended to the next token [sits] for verifying, which outputs [on, a, chair]. By utilizing the hierarchical multi-branch draft, the shared prefix token [on] are merged, after which the token list [my, knee, a, table] are appended. By doing so, given the same decoding length, we are able to save an additional space to append one more token [chair] from another branch [on, a, chair]. Combined with the next token, we collect the token list [sits, on, my, knee, a, table, chair] for decoding through the single forward process and output [on, a, chair, and]. Therefore, compared to the single-branch strategy, our multi-branch strategy, particularly the hierarchical multi-branch draft, offers the advantage of retrieving multiple drafts, resulting in improved EDL from a statistical perspective and thus significantly enhancing LLMs’ inference speed. As illustrated in Figure 3, the position IDs and causal masks in a transformer-based model (Vaswani et al., 2017) are also merged to align with the merged token list. By doing so, we are able to accommodate more branches using hierarchical multi-branch draft, compared to parallel multi-branch draft.
@@ 4.2.2. Hierarchical Multi-Branch Draft
4.3. Trie-tree-based Draft Retrieval
To enable hierarchical multi-branch draft, we utilize a trie tree (Briandais, 1959; Ferragina et al., 2023), a data structure that is widely used for efficient retrieval and storage purposes and handles prefix matching by organizing nodes as individual characters or words. In our work, each node of the trie tree represents a token ID, and a path from the root to a leaf node represents a branch. The trie tree is initialized when a model is loaded, and it is alive until the model instance is shut down.
@@ 4.3.1. Trie Tree Updating
We apply the branch inserting, branch eliminating and node pruning to update the trie tree.
Prompt branch Inserting. Considering that in several typical scenarios (e.g., RAG, Summary), the output are often derived from the prompt, we thus manipulate the prompt into the branches and insert them into the trie tree.
Generated branch Inserting. Different from LLMA which only uses branches from prompts, we also reuse the branch from generated tokens. We find that, towards various prompts, there may be similar outputs, e.g., ”keep exercising” in answering ”how to keep healthy” and ”how to be stronger”. Moreover, we also find that some tokens may be repeated in a response, from example, ”numbers.push_back” may occur several times in a code-generation response. To benefit from the repeat, we put the generated branches into the trie tree on-the-fly instead of the final step.
@@
Node Pruning. To maintain the trie tree within a moderate size, in case that the oversized trie tree results in high memory consumption and slow retrieval performance, we develop a pruning strategy. To achieve this, we decay the branch frequency and remove the nodes with frequency less than 1, when the trie tree exceeds a predetermined capacity. By doing so, we are able to optimize memory consumption and improve retrieval performance in our framework.
@@ 4.3.2. Trie Tree Retrieving
Multi-stage Retrieval. We can extract a sub tree from a trie tree by providing a prefix, which is essentially a list of tokens. The sub tree is also a trie tree and can be directly used for the hierarchical multi-branch draft. The length of the prefix influences the number and relevance of the retrieved branches. Short prefixes yield a greater number of branches, while longer prefixes result in more related branches. To strike a balance between the count and correlation of the branches, we adopt a multi-stage retrieval strategy inspired by (Yang et al., 2023a). Specifically, we begin by attempting to match a longer prefix. If the number of tokens associated with the matched branches is significantly smaller than the CDL, we reduce the length of the prefix and retry the matching process until we obtain a substantial number of tokens linked to matched branches. If the count of the matched branches falls below a predefined threshold, we utilize all of them for the VA process. However, if the count exceeds a given size, we choose the tokens with the highest frequency.
Branch Weighting. For token sorting by frequency, we employ a weighted scheme that considers both prompt and response frequencies. Intuitively, branches from an input prompt may be more related to current generation than that of other responses. To prioritize the importance of branches derived from the input prompt, we amplify their frequency by a significant factor.
5. Experiments
5.1. Experimental Setup
We conduct a solid experiment to evaluate the efficacy, robustness and practicality of our Lookahead framework in accelerating LLMs’ inference. Our inference environment includes a server with 32-core CPU, 64 GB host memory, and various GPUs. Considering the actual industry scenarios that Lookahead is applied, AntRAG is chosen as the evaluation dataset. AntRAG dataset is an internal benchmark dataset that is meticulously collected from a real-life product system in Alipay. The dataset collection process involves submitting user queries to a search engine, which then retrieves the most relevant document alongside the original query. These query-document pairs are combined to form prompts. To ensure the dataset’s quality and coherence, we employ a rigorous process of removing any duplicate or unrelated prompts, resulting in a refined and comprehensive dataset. Furthermore, to validate the robustness and practicality of Lookahead in other scenarios, we additionally conduct experiments utilizing the Dolly dataset (an open domain QA dataset)  [^2] https://huggingface.co/datasets/databricks/databricks-dolly-15k, GSM8k dataset (a dataset of 8.5K high quality linguistically diverse grade school math word problems)  [^3] https://huggingface.co/datasets/gsm8k, and HumanEval-x dataset (a dataset of Python, C++, Java, JavaScript, and Go code tasks)  [^4] https://huggingface.co/datasets/THUDM/HumanEval-x. Detailed information regarding the datasets, base models and devices utilized can be found in the Appendix. For inference speed evaluation, each dataset’s test set is utilized, with the corresponding development set employed for warm-up purposes. It is important to clarify that this warm-up procedure is implemented solely to ensure the convergence and accuracy of performance metrics, as detailed in the Appendix, and is not included in real-world deployment scenarios. In evaluating the candidate methods, which all purport lossless accuracy in generation, inference speed, quantified as output tokens generated per second, serves as the sole metric. Optimal hyper-parameters are determined through a grid search, with the grid size deliberately set as multiples of four to enhance time efficiency.
5.2. Results
5.2.1. Lookahead’s superior performance
Table 2 exhibits the inference speeds achieved by different acceleration methods. The mean value of the inference speed is then selected as the performance indicator for each acceleration method.
Table 2. Inference Speed(token/s) of different methods on various models, datasets and devices. As LookaheadDecoding only offers implementation for Llama, we do not conduct experience on AntGLM-10B Model.
model dataset device transformers vLLM LLMA LaDe lookahead(ours)
AntGLM-10B AntRAG A100-80G 52.4 52.06(x0.99) 165.4(x3.16) - 280.9(x5.36)
AntGLM-10B AntRAG A10 20.3 20.29(x1.00) 59.5(x2.93) - 105.1(x5.18)
AntGLM-10B AntRAG V100-32G 27.3 27.28(x1.00) 64.3(x2.36) - 118.9(x4.36)
Llama-7B Dolly A100-80G 50.4 91.04(x1.81) 60.7(x1.20) 66.7(x1.32) 106.8(x2.12)
Llama-7B Dolly A10 31.4 32.58(x1.04) 35.8(x1.14) 38.2(x1.22) 55.7(x1.77)
Llama-7B GSM8k A100-80G 41.4 92.09(x2.22) 69.0(x1.67) 94.3(x2.28) 111.3(x2.69)
Llama-7B GSM8k A10 31.4 32.73(x1.04) 41.1(x1.31) 48.9(x1.56) 68.1(x2.17)
Llama-7B HumanEval-x A100-80G 51.1 90.47(x1.77) 66.5(x1.30) 88.3(x1.73) 161.5(x3.16)
Llama-7B HumanEval-x A10 30.9 32.46(x1.05) 42.1(x1.36) 46.0(x1.49) 89.6(x2.90)
Llama-13B Dolly A100-80G 39.9 51.67(x1.29) 59.0(x1.48) 45.2(x1.13) 84.6(x2.12)
Llama-13B Dolly V100-32G 20.5 22.07(x1.08) 23.9(x1.17) 23.2(x1.13) 35.2(x1.72)
Llama-13B GSM8k A100-80G 42.9 52.06(x1.21) 51.8(x1.21) 64.8(x1.51) 103.4(x2.41)
Llama-13B GSM8k V100-32G 22.0 22.43(x1.02) 25.8(x1.17) 27.8(x1.26) 45.6(x2.07)
Llama-13B HumanEval-x A100-80G 35.0 51.49(x1.47) 57.7(x1.65) 66.1(x1.89) 137.3(x3.92)
Llama-13B HumanEval-x V100-32G 21.5 22.33(x1.04) 28.8(x1.34) 27.5(x1.28) 57.0(x2.65)
As depicted in Table 2, our Lookahead obtain significant improvement over other methods, on various models, datasets and devices. The average inference speed of AntGLM-10B towards AntRAG is recorded at 52.4 tokens/s on A100 GPU. However, with the incorporation of LLMA, this speed is elevated to 165.4 tokens/s, resulting in a notable 3.16 times speed improvement. Our Lookahead with hierarchical multi-branch draft via trie tree propels the average inference speed even further, culminating in a 5.36 times speed-up, surpassing LLMA by 70% in terms of acceleration. Lookahead also obtains similar acceleration with AntRAG dataset on less-powerful devices, such as A10 and V100, which are widely used for deployment. Despite of RAG scenarios, Lookahead also consistently demonstrates its remarkable superiority and applicability across diverse datasets, especially with speedup of 3.92 on the HumanEval-x dataset, further emphasizing its practicality in real-world scenarios.
5.2.2. Hyper-parameters in multi-branch draft
5.2.3. Hyper-parameters in trie tree updating
Table 3. The inference speed of different updating procedures was evaluated, considering various conditions. In the evaluation, ”W/o prompt” and ”W/o output” refer to the absence of branching from the prompt and output, respectively. ”W/o Pruning” signifies the exclusion of node count reduction in an oversized trie tree through pruning, ”W/o eliminating” denotes the retention of branches derived from the prompt before processing the subsequent query.
Condition W/o prompt W/o output W/o pruning W/o elimin Lookahead
Token/s 234.5 202.0 221.4 234.2 280.9
The capacity of the trie tree has a significant impact on the acceleration performance. To determine the optimal capacity, we conduct empirical experiments, as outlined in Table 4. Instead of using a fixed value, we normalize the node capacity in relation to the decoding length, which allows us to identify a suitable hyper-parameter that performs well across different decoding lengths. Based on our findings, we conclude that node capacities of 16 and 32 times the decoding length yield the best results. Consequently, we have set the default capacity in our framework to be 16 times the decoding length. We also examine the retrieving and updating time of trie tree with different capacity, the time is negligible when compared with the forward time.
Table 4. Inference Speed and retrieving/updating time of different capacity of trie tree. n*DL denotes the capicity is n times the decoding length.
Capacity 1*DL 2*DL 4*DL 8*DL 16*DL 32*DL 64*DL 128*DL
Token/s 254.5 258.7 268.8 273.3 280.7 280.8 279.5 278.3
Retrieve(ms) 0.81 1.35 1.22 1.46 1.02 1.29 1.75 2.38
Update(ms) 0.10 0.12 0.11 0.11 0.12 0.12 0.15 0.16
5.2.4. Lookahead’s efficient memory usage
Figure 5 shows the peak GPU memory consumption for various decoding lengths (1, 2, 4, 8, 16, 32, 64, 128). The decoding length of 1 represents the experience where no Lookahead is used, resulting in a GPU memory consumption of 20.25 GiB. Surprisingly, even with a decoding length of 128, the use of Lookahead only leads to a negligible increase in GPU memory consumption of 0.6%. This minimal increase in memory usage is practically negligible and has no significant impact on real-world applications. Furthermore, we have also evaluated the CPU memory consumption of the trie tree. With the AntRAG dataset, the trie tree only utilizes a mere 260 MiB of memory. This value is negligible when compared to the total memory capacity of a mainstream server.
Table 5. Inference peak memory of different decoding lengths.
Decoding length 1 2 4 8 16 32 64 128
Memory(GiB) 20.25 20.27 20.27 20.27 20.27 20.29 20.38 20.39
5.3. Online Deployment
Our framework has been widely used in real-world scenarios of Alipay since April 2023, due to its accessibility, accuracy-lossless and remarkable acceleration. To be concise, we only report 5 typical scenarios in Table 6. CHART2JSON is a scenario that uses a multi-modal model to convert a chart image to structured content with JSON format, it can achieve extraordinary acceleration due to plenty of template tokens in JSONs. Citizen biz agent, enterprise info QA and health suggestion are RAG scenarios, they are used for answering user questions with reference documents. Medical report summary is a scenario to summarize the content (texts after OCR) of a medical report image. With the assistance of our lookahead framework, the latency of all scenarios is decreased significantly while the generation results are the same as the original step-by-step strategy.
Scenario Baseline Lookahead Speedup
CHART2JSON 12.97 2.07 6.26x
Citizen Biz Agent 1.67 0.32 5.21x
Enterprise Info QA 14.01 2.74 5.11x
Health Suggestion 12.41 2.66 4.66x
Medical Report Summary 3.33 1.25 2.66x
Meanwhile, we have integrated lookahead into other frameworks and obtained additional improvement. vLLM (Kwon et al., 2023) is a widely used framework with state-of-the-art serving throughput, however we may suffer from its huge latency even when only one query is processing. We have implemented lookahead based on vLLM for the single-query situation and obtain 1.6 times acceleration on a real-life scenario about script generation for virtual human, shown in table 7.
Table 7. Inference Latency with Lookahead for vLLM.
Scenario vLLM vLLM+lookahead speedup
Virtual Human Script 5.109 3.203 1.60
6. Conclusion
Appendix A Workflow of lookahead
@@
Appendix B Dataset, Model and Device summary
The experimental setup, including the datasets, models, and devices employed, is delineated in Table 8, Table 9, and Table 10, respectively. Specifically, the Llama2-7b-chat and Llama2-13b-chat models are utilized for the Dolly and GSM8k datasets, while the CodeLlama-7b and CodeLlama-13b models, which are fine-tuned on code dataset with Llama models, are applied to the HumanEval-x dataset. The prepossessed datasets are available in our repository.
@@ Appendix B Dataset, Model and Device summary
Dataset Split #samples #tokens(prompt) #tokens(answer)
@@
Model AntGLM-10B Llama-7B Llama-13B
Params 10.14B 6.74B 12.71B
Vocab size 115328 32000 32000
Layer 48 32 40
Hidden size 4096 4096 5120
Attention head 32 32 40
MLP size 16384 11008 13824
Table 10. Summary of devices.
Device A100-80G A10 V100-32G
Memory 80G HBM2e 24G GDDR6 32G HBM2
Memory bandwidth 2,039GB/s 600GB/s 900GB/s
FLOPs(FP16) 312T 125T 125T
Appendix C Prompt templates.
@@
Similar to (Taori et al., 2023), we construct a prompt for the Dolly dataset with the template in Table 11.
@@
Appendix D Warmup Analysis
To accurately measure inference speed, a warm-up strategy is employed where responses from the development set are preloaded into the trie tree. As illustrated in Figure 6, inference speed is positively correlated with the number of warm-up samples. Given that a model may process hundreds of thousands of requests over its lifespan, it is reasonable to infer that the average speed would exceed that observed during the initial thousands of requests.
Figure 6. Inference speed with various warm-up sample count.
Appendix E Batch Inference
The experiments described were conducted under single-query conditions (i.e., a batch size of 1). It is essential to recognize that batch inference (i.e., batch size ¿ 1) incurs greater computational demand than single-query inference. To ascertain the computational threshold for batch inference, we evaluated the forward time across varying batch sizes, context lengths, and decoding lengths, as depicted in Figure 7. It can be seen that while the batch size is 4 for instance, an increase in decoding length from 5 to 32 does not substantially augment the forward time. Moreover, forward time still remains manageable for a decoding length increasing from 5 to 16 while the batch size is 16. The GPUs still show their redundancies in these scenarios and lookahead in batch inference could still be applied as an effective approach.
Figure 7. GPU Memory of various batch size, context length and decoding length.
To implement lookahead within batch inference parameters, one must navigate variances in KV cache sequence length and attention masks, with implementation specifics available in the repository. To our knowledge, this constitutes the first batch implementation of speculative decoding methodologies. As evidenced in Table 12, lookahead retains its efficacy even with non-unitary batch sizes, demonstrating a notable increase in inference speed relative to traditional step-by-step and LLMA-based batch inference approaches. Despite the less pronounced speedup compared to unitary batch scenarios, two primary factors contribute to this: first, large batch sizes decrease the GPU’s redundancy, particularly with extensive context and decoding lengths, thereby capping potential speed enhancements. Second, the KV cache lengths’ heterogeneity within a batch introduces extra computational overhead in attention operations, diminishing inference velocity.
Nonetheless, it is critical to emphasize that Lookahead is tailored for time-sensitive environments rather than those sensitive to throughput. Future work includes optimizing throughput by incorporating continuous batching (Yu and Jeong, 2022) and high-efficiency attention mechanisms (Dao et al., 2022; Kwon et al., 2023).
Method Batch size=2 Batch size=4
token/s speedup token/s speedup
Baseline 68.0 1.00x 88.2 1.00x
LLMA 185.2 2.72x 214.0 2.43x
Lookahead 285.5 4.20x 299.5 3.40x